# Zhihui Zhu

### v1

@@ 1.1 Background and Motivation
In the past decade, deep learning has exhibited remarkable empirical success across a wide range of engineering and scientific applications [24], such as computer vision [14, 43], natural language processing [48, 49], and health care [11], to name a few. As argued by Bengio et al. [2], Ma et al. [31], one major factor contributing to the success of deep learning is the ability of deep networks to perform powerful nonlinear feature learning by converting the data distribution to a compact and structured representation. This representation greatly facilitates various downstream tasks, including classification [10], segmentation [21], and generation [42].
Based on the theory of data compression and optimal coding [30], Chan et al. [4], Yu et al. [56] proposed a principled and unified framework for deep learning to learn a compact and structured representation. Specifically, they proposed to maximize the difference between the coding rate of all features and the sum of coding rates of features in each class, which is referred to as maximal coding rate reduction (MCR2). This problem is presented in Problem (4) and visualized in Figure 1(a). Here, the coding rate measures the “compactness" of the features, which is interpreted as the volume of a particular set spanned by the learned features: a lower coding rate implies a more compact feature set [^1] Please refer to Chan et al. [4, Section 2.1] for more details on measuring compactness of feature sets via coding rates.. Consequently, the MCR2 objective aims to maximize the volume of the set of all features while minimizing the volumes of the sets of features from each class. Motivated by the structural similarities between deep networks and unrolled optimization schemes for sparse coding [13, 34], Chan et al. [4] constructed a new deep network based on an iterative gradient descent scheme to maximize the MCR2 objective. [^2] When performing maximization, we actually mean that we use gradient ascent. However, we write gradient descent to maintain consistency with existing optimization literature. Notably, each component of this deep network has precise optimization and geometric interpretations. Moreover, it has achieved strong empirical performance on various vision and language tasks [6, 57].
@@
Figure 1: An illustration of the properties of MCR2. (a) The high-dimensional data $\{\bm{x}_{i}\}\subseteq\mathbb{R}^{n}$ lies on a union of low-dimensional submanifolds. The objective of MCR2 is to learn a feature mapping $f_{\bm{\Theta}}(\bm{x})\in\mathbb{R}^{d}$ such that $\bm{z}_{i}=f_{\bm{\Theta}}(\bm{x}_{i})$ for all $i$ are low-dimensional, discriminative, and diverse. (b) According to Theorem 1 and Theorem 2, the regularized MCR2 problem has a benign optimization landscape: each critical point is either a local maximizer or a strict saddle point. Furthermore, each local maximizer, just like the global maximizer, corresponds to a feature representation that consists of a family of orthogonal subspaces, as illustrated in the middle.
@@ 1.2 Our Contributions
Next, we show that the regularized MCR2 function possesses a benign global optimization landscape, despite its complicated structures. More precisely, each critical point is either a local maximizer or strict saddle point of the regularized MCR2 problem; see Figure 1(b). Consequently, any gradient-based optimization, such as (stochastic) gradient descent, with random initialization can escape saddle points and at least converge to a local maximizer efficiently.
@@
Our results not only establish optimization foundations for the MCR2 problem but also yield some important implications for the MCR2-based approach to deep learning. Namely, our theoretical characterizations of local and global optimality offer a compelling explanation for the empirical observations that both deep networks constructed via gradient descent applied to the MCR2 objective and over-parameterized deep networks trained by optimizing the MCR2 objective learn low-dimensional, discriminative, and diverse representations. These results align with the motivations of Chan et al. [4], Yu et al. [56] for employing the MCR2 principle for deep learning, and elucidate the outstanding performance of MCR2-based neural networks across a wide range of vision and language tasks [6, 59]. Moreover, our results underscore the potential of MCR2-based approaches to serve as a cornerstone for future advancements in deep learning, offering a principled approach to pursuing structured and compact representations in practical applications.
@@ 1.3 Related Work
In the literature, it has long been believed that the role of deep networks is to learn certain (nonlinear) low-dimensional and informative representations of the data [16, 31]. For example, Papyan et al. [38] showed that the features learned by cross-entropy (CE) loss exhibit a neural collapse phenomenon during the terminal phase of training, where the features from the same class are mapped to a vector while the features from different classes are maximally linearly separable. Ansuini et al. [1], Recanatesi et al. [40] demonstrated that the dimension of the intermediate features first rapidly increases and then decreases from shallow to deep layers. Masarczyk et al. [32] concluded that the deep layers of neural networks progressively compress within-class features to learn low-dimensional features. Notably, Wang et al. [53] proposed a theoretical framework to analyze hierarchical feature learning for learning low-dimensional representations. They showed that each layer of deep linear networks progressively compresses within-class features and discriminates between-class features in classification problems.
@@
The MCR2-based approach to deep learning for seeking structured and compact representations was first proposed by Yu et al. [56]. Notably, they provided a global optimality analysis of the MCR2 problem (4) with additional rank constraints on the feature matrix of each class. Chan et al. [4] designed a new multi-layer deep network architecture, named ReduNet, based on an iterative gradient descent scheme for maximizing the MCR2 objective. To learn self-consistent representations, Dai et al. [8] extended this approach to the closed-loop transcription (CTRL) framework, which is formulated as a max-min game to optimize a modified MCR2 objective. This game was shown to have global equilibria corresponding to compact and structured representations [37]. Recently, Yu et al. [58] showed that a transformer-like architecture named CRATE, which obtains strong empirical performance [6, 57, 59], can be naturally derived through an iterative optimization scheme for maximizing the sparse rate reduction objective, which is an adaptation to sequence data of the MCR2 objective studied in this work.
@@ 2 Problem Setup
In this section, we first review the basic concepts of MCR2 for deep representation learning in Section 2.1, and then introduce our studied problem in Section 2.2.
@@ 2.1 An Overview of MCR2
In deep representation learning, given data $\{\bm{x}_{i}\}_{i=1}^{m}\subseteq\mathbb{R}^{n}$ from multiple (say $K$) classes, the goal is to learn neural-network representations of these samples that facilitate downstream tasks. Recent empirical studies have shown that good features can be learned for tasks such as classification or autoencoding by using heuristics to promote either the contraction of samples in the same class [41] or the contrast of samples between different classes [15, 36] during the training of neural networks. Notably, Chan et al. [4], Yu et al. [56] unified and formalized these practices and demonstrated that the MCR2 objective is an effective objective to learn within-class compressible and between-class discriminative representations of the data.
@@
On one hand, to make features between different classes discriminative or contrastive, one can maximize the lossy coding rate of all features in $\bm{Z}$, as argued in [4, 56], as follows:
@@
where $\epsilon>0$ is a prescribed quantization error. [^4] Here, $R(\bm{Z})$ is also known as the rate-distortion function in information theory [7], which represents the average number of binary bits needed to encode the data $\bm{Z}$. On the other hand, to make features from the same class compressible or contractive, one can minimize the average lossy coding rate of features in the $k$-th class as follows:
@@
$\max_{\bm{Z}\in\mathbb{R}^{d\times m}}\ R(\bm{Z})-\sum_{k=1}^{K}R_{c}\left(\bm{Z};\bm{\pi}^{k}\right)\quad\mathrm{s.t.}\ \|\bm{Z}_{k}\|_{F}^{2}=m_{k},\quad\forall k\in[K].$ (4)
This is referred to as the principle of maximal coding rate reduction in [4, 56]. It is worth mentioning that this principle can be extended to self-supervised and even unsupervised learning settings, where we learn the label vectors $\{\bm{\pi}^{k}\}_{k=1}^{K}$ during training.
@@ 2.2 The Regularized MCR2 Problem
Due to the Frobenius norm constraints, it is a tremendously difficult task to analyze Problem (4) from an optimization-theoretic perspective, as all the analysis would occur on a product of spheres instead of on Euclidean space. Therefore, we consider the Lagrangian formulation of (4). This can be viewed as a tight relaxation or even an equivalent problem of (4) whose optimal solutions agree under specific settings of the regularization parameter; see Proposition 1. Specifically, the formulation we study, referred to henceforth as the regularized MCR2 problem, is as follows:
@@
The first approach, as argued by Chan et al. [4], Yu et al. [57], is to construct a new deep network architecture, i.e., ReduNet [4] or CRATE [57], based on an iterative gradient descent scheme to optimize the MCR2-type objective. In this approach, each layer of the constructed network approximates a gradient descent step to optimize the MCR2-type objective given the input representation. The key takeaway is that these networks approximately implement gradient descent directly on the representations, so our analysis of the optimization properties of the MCR2-type objective translates to explanations of the corresponding properties of the learned representations and architectures of these deep networks. In particular, our argument that the optima and optimization landscape of (5) are favorable directly translates to a justification of the correctness of learned representations of the ReduNet and a validation of its architecture design. Moreover, this study enables principled improvement of deep network architectures constructed via unrolled optimization by leveraging more advanced optimization techniques better suited for problems with benign landscapes. This can improve model interpretability and efficiency.
@@
In the second approach, one parameterizes the feature mapping $f_{\bm{\Theta}}(\cdot)$ via standard deep neural networks such as a multi-layer perceptron or ResNet [14], and treats the MCR2-type objective like other loss functions applied to outputs of a neural network, such as mean-squared error or cross-entropy loss. Studying Problem (5) from this perspective would require us to optimize over $\bm{\Theta}$ instead of over $\bm{Z}$. This new optimization problem would be extraordinarily difficult to analyze, because modern neural networks have nonlinear interactions across many layers, so the parameters $\bm{\Theta}$ would affect the final representation $\bm{Z}$ in a complex way. Fortunately, since modern neural networks are often highly over-parameterized, they can interpolate or approximate any continuous function in the feature space [29], so we may omit these constraints by assuming the unconstrained feature model, where $\bm{z}_{i}$ for all $i\in[N]$ are treated as free optimization variables [33, 55, 62, 51]. Consequently, studying the optimization properties of Problem (5) provides valuable insights into the structures of learned representations and the efficiency of training deep networks using MCR2-type objectives.
@@ 3 Main Results
In this section, we first characterize the local and global optimal solutions of Problem (5) in Section 3.1, and then analyze the global landscape of the objective function in Section 3.2.
@@ 3.1 Characterization of Local and Global Optimality
We defer the proof to Section 4.1 and Section D.1. In this theorem, we explicitly characterize the local and global optima of Problem (5). Intuitively, this demonstrates that the features represented by each local maximizer of Problem (5) are low-dimensional and discriminative in the sense that
@@
Our above discussion explains the merits of achieving both local and global optima. At each maximizer, the representations are within-class compressible and between-class discriminative (Theorem 1 (i)). Moreover, global maximizers further satisfy that the representations are all maximally diverse (Theorem 1 (ii)(a)). If all classes were balanced, i.e., $m_{1}=\cdots=m_{K}$, then Theorem 1 (ii)(b) would not apply, and these properties would be all that Theorem 1 asserts. In this case, global optima would clearly be desired over local optima. However, in the unbalanced case, the situation is more complex, because Theorem 1 (ii)(b) would apply. It says that for global optima, the classes with the smallest numbers of samples would fill to the largest dimension possible, and the very largest classes could collapse to $\bm{0}$, an undesirable situation. A dramatic example of this is when $m_{1}>\cdots>m_{K}>d$, for then any global optimum would have $\operatorname{rank}(\bm{Z}_{K})=d$ and $\bm{Z}_{1},\dots,\bm{Z}_{K-1}$ all collapse to $\bm{0}$. Overall, in the unbalanced case, global optima may not always correspond to the best representations. In particular, local optima with more equitable rank distributions (like bigger classes span more dimensions) which are still maximally diverse (i.e., ranks of each class sum to the dimension $d$) could be preferred in applications. As demonstrated in Section 5.1, these kinds of potentially useful local optima are realized in experiments, even with unbalanced classes.
@@
Based on the characterization of global optimality in Theorem 1, we show the following proposition that establishes the relationship between the constrained MCR2 problem (4) and the regularized MCR2 problem (5) in terms of their global solutions under an appropriate choice of the regularization parameter. The proof of this result can be found in Section D.2.
@@ 3.2 Analysis of Global Optimization Landscape
While we have characterized the local and global optimal solutions in Theorem 1, it remains unknown whether these solutions can be computed efficiently using GD to solve Problem (5), as GD may get stuck at a saddle point. Fortunately, Sun et al. [44], Lee et al. [25] showed that if a function is twice continuously differentiable and satisfies strict saddle property, i.e., each critical point is either a local minimizer or a strict saddle point [^5] We say that a critical point is a strict saddle point of Problem (5) if it has a direction with strictly positive curvature; see Definition 2. This includes classical saddle points with strictly positive curvature as well as local minimizers., GD converges to its local minimizer almost surely with random initialization. We investigate the global optimization landscape of Problem (5) by characterizing all of its critical points as follows.
@@
We defer the proof to Section 4.2 and Section D.3. Here, we make some remarks on this theorem and also on the consequences of the results derived so far.
@@
Chan et al. [4], Yu et al. [56] have characterized the global optimality of Problem (4) with Frobenius norm constraints on each $\bm{Z}_{k}$ in the UFM. However, their analysis requires an additional rank constraint on each $\bm{Z}_{k}$ and only characterizes globally optimal representations. In contrast, our analysis eliminates the need for the rank constraint, and we characterize local and global optimality in Problem (5), as well as its optimization landscape. Interestingly, we demonstrate that the features represented by each local maximizer — not just global maximizers — are also compact and structured. Furthermore, we demonstrate that the regularized MCR2 objective (5) is a strict saddle function. To the best of our knowledge, Theorems 1 and 2 constitute the first analysis of local optima and optimization landscapes for MCR2 objectives. According to Daneshmand et al. [9], Lee et al. [25], Xu et al. [54], Theorems 1 and 2 imply that low-dimensional and discriminative representations can be efficiently found by (stochastic) GD on Problem (5) from a random initialization.
@@
In recent years, there has been a growing body of literature exploring optimization landscapes of non-convex problems in machine learning and deep learning. These include low-rank matrix factorization [12, 46, 5, 60], community detection [50, 52], dictionary learning [45, 39], and deep neural networks [47, 55, 61, 62, 18, 27]. The existing analyses in the literature cannot be applied to the MCR2 problem due to its special structure, which involves the log-determinant of all features minus the sum of the log-determinant of features in each class. Our work contributes to the literature on optimization landscape analyses of non-convex problems by showing that the MCR2 problem has a benign optimization landscape. Our approach may be of interest to analyses of the landscapes of other intricate loss functions in practical applications.

### v2

@@ 1.1 Background and Motivation
In the past decade, deep learning has exhibited remarkable empirical success across a wide range of engineering and scientific applications [lecun2015deep], such as computer vision [he2016deep, simonyan2014very], natural language processing [sutskever2014sequence, vaswani2017attention], and health care [esteva2019guide], to name a few. As argued by bengio2013representation, ma2022principles, one major factor contributing to the success of deep learning is the ability of deep networks to perform powerful nonlinear feature learning by converting the data distribution to a compact and structured representation. This representation greatly facilitates various downstream tasks, including classification [dosovitskiy2020image], segmentation [kirillov2023segment], and generation [saharia2022photorealistic].
Based on the theory of data compression and optimal coding [ma2007segmentation], chan2022redunet, yu2020learning proposed a principled and unified framework for deep learning to learn a compact and structured representation. Specifically, they proposed to maximize the difference between the coding rate of all features and the sum of coding rates of features in each class, which is referred to as maximal coding rate reduction (MCR2). This problem is presented in Problem (4) and visualized in Figure˜1(a). Here, the coding rate measures the “compactness" of the features, which is interpreted as the volume of a particular set spanned by the learned features: a lower coding rate implies a more compact feature set [^1] Please refer to chan2022redunet for more details on measuring compactness of feature sets via coding rates.. Consequently, the MCR2 objective aims to maximize the volume of the set of all features while minimizing the volumes of the sets of features from each class. Motivated by the structural similarities between deep networks and unrolled optimization schemes for sparse coding [gregor2010learning, monga2021algorithm], chan2022redunet constructed a new deep network based on an iterative gradient descent scheme to maximize the MCR2 objective. [^2] When performing maximization, we actually mean that we use gradient ascent. However, we write gradient descent to maintain consistency with existing optimization literature. Notably, each component of this deep network has precise optimization and geometric interpretations. Moreover, it has achieved strong empirical performance on various vision and language tasks [chu2023image, yu2023sparse].
@@
Figure 1: An illustration of the properties of MCR2. (a) The high-dimensional data $\{\bm{x}_{i}\}\subseteq\mathbb{R}^{n}$ lies on a union of low-dimensional submanifolds. The objective of MCR2 is to learn a feature mapping $f_{\bm{\Theta}}(\bm{x})\in\mathbb{R}^{d}$ such that $\bm{z}_{i}=f_{\bm{\Theta}}(\bm{x}_{i})$ for all $i$ are low-dimensional, discriminative, and diverse. (b) According to Theorem˜1 and Theorem˜2, the regularized MCR2 problem has a benign optimization landscape: each critical point is either a local maximizer or a strict saddle point. Furthermore, each local maximizer, just like the global maximizer, corresponds to a feature representation that consists of a family of orthogonal subspaces, as illustrated in the middle.
@@ 1.2 Our Contributions
Next, we show that the regularized MCR2 function possesses a benign global optimization landscape, despite its complicated structures. More precisely, each critical point is either a local maximizer or strict saddle point of the regularized MCR2 problem; see Figure˜1(b). Consequently, any gradient-based optimization, such as (stochastic) gradient descent, with random initialization can escape saddle points and at least converge to a local maximizer efficiently.
@@
Our results not only establish optimization foundations for the MCR2 problem but also yield some important implications for the MCR2-based approach to deep learning. Namely, our theoretical characterizations of local and global optimality offer a compelling explanation for the empirical observations that both deep networks constructed via gradient descent applied to the MCR2 objective and over-parameterized deep networks trained by optimizing the MCR2 objective learn low-dimensional, discriminative, and diverse representations. These results align with the motivations of chan2022redunet, yu2020learning for employing the MCR2 principle for deep learning, and elucidate the outstanding performance of MCR2-based neural networks across a wide range of vision and language tasks [chu2023image, yu2023emergence]. Moreover, our results underscore the potential of MCR2-based approaches to serve as a cornerstone for future advancements in deep learning, offering a principled approach to pursuing structured and compact representations in practical applications.
@@ 1.3 Related Work
In the literature, it has long been believed that the role of deep networks is to learn certain (nonlinear) low-dimensional and informative representations of the data [hinton2006reducing, ma2022principles]. For example, papyan2020prevalence showed that the features learned by cross-entropy (CE) loss exhibit a neural collapse phenomenon during the terminal phase of training, where the features from the same class are mapped to a vector while the features from different classes are maximally linearly separable. ansuini2019intrinsic, recanatesi2019dimensionality demonstrated that the dimension of the intermediate features first rapidly increases and then decreases from shallow to deep layers. masarczyk2023tunnel concluded that the deep layers of neural networks progressively compress within-class features to learn low-dimensional features. Notably, wang2023understanding proposed a theoretical framework to analyze hierarchical feature learning for learning low-dimensional representations. They showed that each layer of deep linear networks progressively compresses within-class features and discriminates between-class features in classification problems.
@@
The MCR2-based approach to deep learning for seeking structured and compact representations was first proposed by yu2020learning. Notably, they provided a global optimality analysis of the MCR2 problem (4) with additional rank constraints on the feature matrix of each class. chan2022redunet designed a new multi-layer deep network architecture, named ReduNet, based on an iterative gradient descent scheme for maximizing the MCR2 objective. To learn self-consistent representations, dai2022ctrl extended this approach to the closed-loop transcription (CTRL) framework, which is formulated as a max-min game to optimize a modified MCR2 objective. This game was shown to have global equilibria corresponding to compact and structured representations [pai2022pursuit]. Recently, yu2023white showed that a transformer-like architecture named CRATE, which obtains strong empirical performance [chu2023image, yu2023sparse, yu2023emergence], can be naturally derived through an iterative optimization scheme for maximizing the sparse rate reduction objective, which is an adaptation to sequence data of the MCR2 objective studied in this work.
@@ 2 Problem Setup
In this section, we first review the basic concepts of MCR2 for deep representation learning in Section˜2.1, and then introduce our studied problem in Section˜2.2.
@@ 2.1 An Overview of MCR2
In deep representation learning, given data $\{\bm{x}_{i}\}_{i=1}^{m}\subseteq\mathbb{R}^{n}$ from multiple (say $K$) classes, the goal is to learn neural-network representations of these samples that facilitate downstream tasks. Recent empirical studies have shown that good features can be learned for tasks such as classification or autoencoding by using heuristics to promote either the contraction of samples in the same class [rifai2011contractive] or the contrast of samples between different classes [he2020momentum, oord2018representation] during the training of neural networks. Notably, chan2022redunet, yu2020learning unified and formalized these practices and demonstrated that the MCR2 objective is an effective objective to learn within-class compressible and between-class discriminative representations of the data.
@@
On one hand, to make features between different classes discriminative or contrastive, one can maximize the lossy coding rate of all features in $\bm{Z}$, as argued in [chan2022redunet, yu2020learning], as follows:
@@
where $\epsilon>0$ is a prescribed quantization error. [^4] Here, $R(\bm{Z})$ is also known as the rate-distortion function in information theory [cover1999elements], which represents the average number of binary bits needed to encode the data $\bm{Z}$. On the other hand, to make features from the same class compressible or contractive, one can minimize the average lossy coding rate of features in the $k$-th class as follows:
@@
$\max_{\bm{Z}\in\mathbb{R}^{d\times m}}\ R(\bm{Z})-\sum_{k=1}^{K}R_{c}\left(\bm{Z};\bm{\pi}^{k}\right)\qquad\mathrm{s.t.}\ \|\bm{Z}_{k}\|_{F}^{2}=m_{k},\qquad\forall k\in[K].$ (4)
This is referred to as the principle of maximal coding rate reduction in [chan2022redunet, yu2020learning]. It is worth mentioning that this principle can be extended to self-supervised and even unsupervised learning settings, where we learn the label vectors $\{\bm{\pi}^{k}\}_{k=1}^{K}$ during training.
@@ 2.2 The Regularized MCR2 Problem
Due to the Frobenius norm constraints, it is a tremendously difficult task to analyze Problem (4) from an optimization-theoretic perspective, as all the analysis would occur on a product of spheres instead of on Euclidean space. Therefore, we consider the Lagrangian formulation of (4). This can be viewed as a tight relaxation or even an equivalent problem of (4) whose optimal solutions agree under specific settings of the regularization parameter; see ˜1. Specifically, the formulation we study, referred to henceforth as the regularized MCR2 problem, is as follows:
@@
The first approach, as argued by chan2022redunet, yu2023sparse, is to construct a new deep network architecture, i.e., ReduNet [chan2022redunet] or CRATE [yu2023sparse], based on an iterative gradient descent scheme to optimize the MCR2-type objective. In this approach, each layer of the constructed network approximates a gradient descent step to optimize the MCR2-type objective given the input representation. The key takeaway is that these networks approximately implement gradient descent directly on the representations, so our analysis of the optimization properties of the MCR2-type objective translates to explanations of the corresponding properties of the learned representations and architectures of these deep networks. In particular, our argument that the optima and optimization landscape of (5) are favorable directly translates to a justification of the correctness of learned representations of the ReduNet and a validation of its architecture design. Moreover, this study enables principled improvement of deep network architectures constructed via unrolled optimization by leveraging more advanced optimization techniques better suited for problems with benign landscapes. This can improve model interpretability and efficiency.
@@
In the second approach, one parameterizes the feature mapping $f_{\bm{\Theta}}(\cdot)$ via standard deep neural networks such as a multi-layer perceptron or ResNet [he2016deep], and treats the MCR2-type objective like other loss functions applied to outputs of a neural network, such as mean-squared error or cross-entropy loss. Studying Problem (5) from this perspective would require us to optimize over $\bm{\Theta}$ instead of over $\bm{Z}$. This new optimization problem would be extraordinarily difficult to analyze, because modern neural networks have nonlinear interactions across many layers, so the parameters $\bm{\Theta}$ would affect the final representation $\bm{Z}$ in a complex way. Fortunately, since modern neural networks are often highly over-parameterized, they can interpolate or approximate any continuous function in the feature space [lu2017expressive], so we may omit these constraints by assuming the unconstrained feature model, where $\bm{z}_{i}$ for all $i\in[N]$ are treated as free optimization variables [mixon2020neural, yaras2022neural, zhu2021geometric, wang2022linear]. Consequently, studying the optimization properties of Problem (5) provides valuable insights into the structures of learned representations and the efficiency of training deep networks using MCR2-type objectives.
@@ 3 Main Results
In this section, we first characterize the local and global optimal solutions of Problem (5) in Section˜3.1, and then analyze the global landscape of the objective function in Section˜3.2.
@@ 3.1 Characterization of Local and Global Optimality
We defer the proof to Section˜4.1 and Section˜D.1. In this theorem, we explicitly characterize the local and global optima of Problem (5). Intuitively, this demonstrates that the features represented by each local maximizer of Problem (5) are low-dimensional and discriminative in the sense that
@@
Our above discussion explains the merits of achieving both local and global optima. At each maximizer, the representations are within-class compressible and between-class discriminative (Theorem˜1 (i)). Moreover, global maximizers further satisfy that the representations are all maximally diverse (Theorem˜1 (ii)(a)). If all classes were balanced, i.e., $m_{1}=\cdots=m_{K}$, then Theorem˜1 (ii)(b) would not apply, and these properties would be all that Theorem˜1 asserts. In this case, global optima would clearly be desired over local optima. However, in the unbalanced case, the situation is more complex, because Theorem˜1 (ii)(b) would apply. It says that for global optima, the classes with the smallest numbers of samples would fill to the largest dimension possible, and the very largest classes could collapse to $\bm{0}$, an undesirable situation. A dramatic example of this is when $m_{1}>\cdots>m_{K}>d$, for then any global optimum would have $\operatorname{rank}(\bm{Z}_{K})=d$ and $\bm{Z}_{1},\dots,\bm{Z}_{K-1}$ all collapse to $\bm{0}$. Overall, in the unbalanced case, global optima may not always correspond to the best representations. In particular, local optima with more equitable rank distributions (like bigger classes span more dimensions) which are still maximally diverse (i.e., ranks of each class sum to the dimension $d$) could be preferred in applications. As demonstrated in Section˜5.1, these kinds of potentially useful local optima are realized in experiments, even with unbalanced classes.
@@
Based on the characterization of global optimality in Theorem˜1, we show the following proposition that establishes the relationship between the constrained MCR2 problem (4) and the regularized MCR2 problem (5) in terms of their global solutions under an appropriate choice of the regularization parameter. The proof of this result can be found in Section˜D.2.
@@ 3.2 Analysis of Global Optimization Landscape
While we have characterized the local and global optimal solutions in Theorem˜1, it remains unknown whether these solutions can be computed efficiently using GD to solve Problem (5), as GD may get stuck at a saddle point. Fortunately, sun2015nonconvex, lee2016gradient showed that if a function is twice continuously differentiable and satisfies strict saddle property, i.e., each critical point is either a local minimizer or a strict saddle point [^5] We say that a critical point is a strict saddle point of Problem (5) if it has a direction with strictly positive curvature; see Definition 2. This includes classical saddle points with strictly positive curvature as well as local minimizers., GD converges to its local minimizer almost surely with random initialization. We investigate the global optimization landscape of Problem (5) by characterizing all of its critical points as follows.
@@
We defer the proof to Section˜4.2 and Section˜D.3. Here, we make some remarks on this theorem and also on the consequences of the results derived so far.
@@
chan2022redunet, yu2020learning have characterized the global optimality of Problem (4) with Frobenius norm constraints on each $\bm{Z}_{k}$ in the UFM. However, their analysis requires an additional rank constraint on each $\bm{Z}_{k}$ and only characterizes globally optimal representations. In contrast, our analysis eliminates the need for the rank constraint, and we characterize local and global optimality in Problem (5), as well as its optimization landscape. Interestingly, we demonstrate that the features represented by each local maximizer — not just global maximizers — are also compact and structured. Furthermore, we demonstrate that the regularized MCR2 objective (5) is a strict saddle function. To the best of our knowledge, Theorems˜1 and 2 constitute the first analysis of local optima and optimization landscapes for MCR2 objectives. According to daneshmand2018escaping, lee2016gradient, xu2018first, Theorems˜1 and 2 imply that low-dimensional and discriminative representations can be efficiently found by (stochastic) GD on Problem (5) from a random initialization.
@@
In recent years, there has been a growing body of literature exploring optimization landscapes of non-convex problems in machine learning and deep learning. These include low-rank matrix factorization [ge2017no, sun2018geometric, chi2019nonconvex, zhang2020symmetry], community detection [wang2021optimal, wang2022non], dictionary learning [sun2017complete_a, qu2020geometric], and deep neural networks [sun2020global, yaras2022neural, zhou2022optimization, zhu2021geometric, jiang2024generalized, li2024neural]. The existing analyses in the literature cannot be applied to the MCR2 problem due to its special structure, which involves the log-determinant of all features minus the sum of the log-determinant of features in each class. Our work contributes to the literature on optimization landscape analyses of non-convex problems by showing that the MCR2 problem has a benign optimization landscape. Our approach may be of interest to analyses of the landscapes of other intricate loss functions in practical applications.